OpenStudy (anonymous):

Are f(x)=(2)/(3)x+6 and g(x)=(3)/(2)x-9 inverse of each other

OpenStudy (4n1m0s1ty):

An inverse function is one that satisfies f(g(x))=x

OpenStudy (4n1m0s1ty):

So try picking a number and plugging it into g(x), then take the result and replace it for x in the f(x) function

OpenStudy (anonymous):

they are not?

OpenStudy (4n1m0s1ty):

So lets say x =2, then g(2) = (3/2)*2 - 9 = -6 then f(g(2)) = f(-6) = (2/3)*-6 + 6 = 2

OpenStudy (4n1m0s1ty):

We started with x = 2 and ended with 2 as the result, so they are inverses of each other